OpenStudy (anonymous):

Riley is saving up to buy a new television. She needs a total of $1,039.00. Riley has saved $200.03 already and earns $76.27 per week at her job. How many weeks does Riley need to work to save enough money to buy the new television? 11 12 10 9

OpenStudy (anonymous):

@asnaseer

Mehek (mehek14):

this is your function: 1039 = 76.27x + 200.03

OpenStudy (asnaseer):

This is very similar to last question. So first subtract $200.03 from $1039.00

OpenStudy (anonymous):

838.97 @asn\

OpenStudy (asnaseer):

good - now you knowRiley earns $76.27 per week. so you just need to divide 838.97 by 76.97 to get the number of weeks required to work

OpenStudy (anonymous):

10.89

OpenStudy (anonymous):

@asnaseer

OpenStudy (asnaseer):

Sorry - my mistake - I meant to say - remember you need to divide 838.97 by 76.27

OpenStudy (asnaseer):

not 76.97

OpenStudy (anonymous):

11

OpenStudy (anonymous):

At their going-out-of-business sale, Bill's Bikes sold 33 bikes and had 59 left. Which equation could be used to find x, the number of bikes the store had originally? x + 59 = -33 x - 33 = -59 x + 33 = 59 x - 59 = 33

OpenStudy (asnaseer):

11 is correct - well done again! :) Please post each question separately in the list on the left - that helps others find similar questions that they may be interested more easily.